OpenStudy (anonymous):

the sum of the ages of tony and teddy equals thrice their mothers age. Tony is two years younger than teddy. if the age of the mother is 42, find the ages of tony and teddy?

OpenStudy (anonymous):

tony and tedy equals thrice their mom age , therefore :y+x=3z , tony is two years younger than teddy :x-y=2 from the first equation : y+x=3z , y+x=3(42), y+x=126 from the second equation : x-y=2 ,x=2+y. subs equ 2 in equ 1 = y+2+y=126 , 2y+2=126 ,2y=124, y=62. subs in equ 2 : x-y=2 ,x-62=2 ,x=64. subs in equ 1 , y+x=3z ,62+64=3z, 126=3z , z =42 . _________________________________________________________ where x=tedy's age , y=tony age , z=moms age :)

OpenStudy (anonymous):

did u understand it hun ? or need more explanation ?